OpenStudy (anonymous):

suppose ln2=a and ln3=b. Use properties of logs to write each in terms of a and b. ln(24)

OpenStudy (anonymous):

can you please include all steps! thank you!!!

pooja195 (pooja195):

ln1.5 = ln(3/2) = ln3 - ln2 = b - a

OpenStudy (anonymous):

thank you!! is ln1.5 the final answer? What does the ln(24) mean?

pooja195 (pooja195):

yes

OpenStudy (anonymous):

ok thanks so much